ERRATA

J. B. Dence, “Primitive Roots the Cyclotomic Way,” Missouri Journal of Mathematical Sciences, 12 (2000), 5–11.

In the 4th column of Table 3 the entries −1, −1, 1, 1 should read 1, 1, −1, −1. Also, in line 3 of page 7, the expression

∏_{d_{i}|d} Φ_{d}(x)

should read

∏_{d_{i}|d} Φ_{d_{i}}(x).